# Scalable Multi-Objective and Meta Reinforcement Learning via Gradient Estimation

### v1

@@ Abstract
Introduction
Reinforcement learning (RL) is a technique for sequential decision-making in interactive environments, enabling agents to learn from feedback signals. A central challenge in RL is to develop agents that generalize across a wide range of tasks rather than solving a single task in isolation. For instance, a robot needs to master multiple related skills that share an underlying structure (yu2020metaworld). Prior work has studied this problem for two closely related settings. The first, multitask reinforcement learning, seeks to find a single policy that performs well across a given set of objectives (or tasks) (sodhani2021multi; joshi2025benchmarking). The second, meta-reinforcement learning, aims to learn a meta-initialization that can rapidly adapt to unseen tasks (finn2017model; rakelly2019efficient). In both settings, however, the computational cost of evaluating and adapting a shared policy over many competing objectives grows quickly with $n$.
@@ Introduction
This paper studies the problem of designing a (scalable) multi-objective optimizer for reinforcement learning, with extensions to policy meta-adaptation. Learning from a diverse set of tasks has been shown to improve generalization (thrun1998learning; cobbe2019quantifying; yin2019meta). However, the main difficulty is modeling task relationships and how different tasks transfer information to one another when trained together in a complex network (wuunderstanding). Existing approaches have considered measuring the similarity between gradient vectors across tasks to quantify task relationships (yu2020pcgrad). Notably, this gradient-similarity-based measure corresponds to pairwise task-affinity scores, whereas in practice it is natural to model how multiple policies interact when trained together (li2023identification). Exhaustively searching over all subsets of $\left\{1,2,\dots,n\right\}$ is computationally infeasible: evaluating each subset typically requires a full training procedure, and the number of subsets grows exponentially in $n$ (caruana1997multitask). Greedy stepwise selection requires evaluation on $O(n^{2})$ subsets, which is again impractical for large $n$ (li2023boosting; li2024scalable).
@@
To overcome these challenges, we introduce a scalable algorithm for estimating policy performance on arbitrary task subsets without full training. Our algorithm starts by learning a single meta-policy across all tasks, then uses a first-order surrogate model—derived from a Taylor approximation around the meta-policy to efficiently estimate the outcome of fine-tuning on any subset. This approach is analogous to a random ensemble (zhang2025linear), which allows us to compute a pairwise task affinity matrix and partition the $n$ objectives into $k$ subgroups using a convex clustering procedure. The resulting groups can be trained separately using any multitask or meta-RL optimizers. This overall approach is illustrated in Figure 1.
@@
We validate our approach through extensive experiments on both Meta-World (yu2020metaworld) and robotic control benchmarks (towers2024gymnasium). We find that the first-order approximation applied to the outputs of a policy network yields accurate estimates—typically within $2\%$ error margin—while being up to $26\times$ faster than full training. Our algorithm consistently outperforms multitask and meta-RL baselines (yang2020multi; sodhani2021multi; sun2022paco) by $16\%$ on average, and surpasses baseline grouping strategies (such as random grouping or gradient-similarity based grouping) by $23\%$ and $16\%$, respectively. These results highlight the effectiveness and scalability of our task-affinity estimation approach.
@@
Lastly, we conduct a theoretical analysis of the generalization errors of policy networks. The main technical approach is to compute (and measure) the trace of the Hessian of the loss surface (ju2022robust; ju2023generalization; zhangnoise). See Theorem 1 for the precise statement. When evaluated on the above RL environments, we find that the Hessian-based generalization bounds match the scale of empirical generalization errors (see Figure 3).
@@
Pairwise task interactions can be characterized through this notion. For example, $R_{\{i,j\}}(\theta)$ captures the joint performance on tasks $T_{i}$ and $T_{j}$, providing a meaningful measure of their affinity. A key challenge in multitask learning is negative transfer, where conflicting gradients from two tasks degrade optimization efficiency and final performance (wuunderstanding). This occurs when
@@
Our goal is to model the underlying relationship between the $n$ tasks, which can be used to group similar tasks or select a small representative subset for meta-learning. However, searching for the optimal partition requires evaluating $R_{S}(\theta)$ for multiple subsets $S\subseteq\mathcal{T}$, since the number of possible subsets is $2^{n}$. Greedy stepwise selection again requires evaluating $\mathcal{O}(n^{2})$ subsets, where each evaluation typically involves a full RL training or adaptation procedure (li2023boosting; li2024scalableft). This is again intractable for large $n$.
@@
We begin with the standard policy gradient objective used in algorithms such as Proximal Policy Optimization (PPO) (schulman2017proximal):
@@
Random projection. Since the gradient vectors $g_{i,t}\in\mathbb{R}^{p}$ may be high-dimensional (with $p$ in the order of millions), we apply random projections to reduce the dimension down to a few hundred (in practice, $d=400$ suffices). This projection provably preserves the pairwise similarity between all pairs of gradient vectors via the Johnson–Lindenstrauss Lemma (johnson1984extensions). We sample a random projection matrix $P\in\mathbb{R}^{p\times d}$, where $d\ll p$ and every entry of $P$ is independently drawn from a Gaussian $\mathcal{N}(0,d^{-1})$. We project the high-dimensional gradients to a $d$-dimensional space as
@@
Evaluation of First-Order Policy Approximation
@@
Task Affinity Grouping
@@
We then cluster tasks by solving a convex optimization problem that maximizes intra-cluster affinity, following a trace-regularized relaxation (awasthi2015relax). This step is very fast since it runs on an $n$ by $n$ matrix with $n$ being at most several hundred, which can typically be solved in just a few seconds. Moreover, this procedure can be repeated with different values of $\lambda$ to determine the optimal number of clusters $k$ for downstream performance. For brevity, we defer a complete description of the convex relaxation program to the Appendix. The whole procedure is described in Algorithm 2.
@@
RL Environment Meta-World CartPole Highway LunarLander
Multi-task training (yu2020metaworld) $71.3_{\pm 1.2}\%$ $145.9_{\pm 9.0}$ $140.0_{\pm 4.6}$ $53.8_{\pm 14.6}$
Soft modularization (yang2020multi) $82.0_{\pm 1.1}\%$ $139.3_{\pm 9.5}$ $141.3_{\pm 3.5}$ $66.1_{\pm 13.0}$
PaCo (sun2022paco) $73.1_{\pm 1.1}\%$ $144.5_{\pm 5.2}$ $136.6_{\pm 6.4}$ $62.6_{\pm 11.4}$
CARE (sodhani2021multi) $84.0_{\pm 1.8}\%$ / / /
Gradient-similarity-based grouping (yu2020pcgrad) $69.6_{\pm 1.9}\%$ $142.0_{\pm 8.2}$ $135.6_{\pm 7.6}$ $80.8_{\pm 6.9}$
Table 2: Comparison of our approach with several baseline methods. We report the average success rate on the Meta-World benchmark and the rewards (after adaptation) for three robotic control environments. We note that CARE does not apply to the control tasks in the meta-RL setting. We report the mean and standard deviation from five runs.
Experiments
@@
Experimental Setup
Environments. We evaluate our method on two types of benchmarks. First, we use MT10 from the Meta-World benchmark (yu2020metaworld), which consists of 10 diverse robotic manipulation tasks. While these tasks share common state and action spaces, their reward functions and dynamics vary (e.g., a goal position in one task may correspond to an object’s position in another). Second, we use three classic control environments from Gymnasium (towers2024gymnasium) and MO-Gymnasium (felten_toolkit_2023): CartPole, Highway, and LunarLander. For each, we generate a distribution of 10 source tasks by altering key physical parameters (e.g., pole length for CartPole, traffic density for Highway, and gravity for LunarLander).
Baselines. We compare our method against two categories of baselines. (1) Multi-objective RL baselines: These represent standard and state-of-the-art multitask approaches for multitasks training: a single policy trained on all tasks; Soft Modularization (yang2020multi), which employs a routing-based selection mechanism; PaCo (sun2022paco), an ensemble method that composes policies from a shared parameter subspace; and CARE (sodhani2021multi), which performs inference using contextual information to adapt policies. (2) Grouping baselines: These methods use the same number of groups as our approach but employ different heuristic grouping procedures. This allows us to isolate the benefit of our affinity modeling. We include random grouping (where tasks are randomly assigned to $k$ groups) and gradient-similarity-based grouping (where tasks are clustered based on the cosine similarity of their respective policy gradients).
@@
For the control environments, we form the meta-training set by randomly selecting one from each task group provided by PolicyGradEx. We then use MAML (finn2017model) to train a meta-policy on the assigned tasks.
@@
Results for Loss Function Estimation
@@
(a) Illustration of Hessian traces and generalization errors during training
@@
Results for Multi-Objective RL
Generalization Error Measurements via Hessians
Next, we analyze the generalization behavior. by evaluating a sharpness measure based on the Hessian trace of the test loss for the policy network and the policy gradient loss. A smaller trace value suggests a flatter loss landscape. We implement Hutchinson’s estimator and leverage a faster version of this estimator (meyer2021hutch++) to estimate the Hessian trace of RL models. We compute the policy gradient loss from the target tasks and measure the difference between the training loss and test loss as the generalization error. Results for the largest eigenvalue of the Hessian are similar.
@@
Generalization Error Analysis via Hessians
@@
The proof of Theorem 1 is based on a PAC-Bayes analysis (ju2022robust; zhangnoise), and the details are in the Appendix. In particular, a new aspect of this result is that we use an anisotropic perturbation in the prior and posterior distributions. We build on a linear PAC-Bayes bound (See Theorem 2 in the Appendix), and optimize the noise distribution to obtain the $\mathcal{H}$ measure.
@@
Related Work
Multi-objective optimization for reinforcement learning. Early works on multitask learning use hard parameter sharing, where a single neural network backbone is shared among all tasks, with only the final layers or heads being task-specific (yu2020metaworld). A canonical example is the multi-head Soft Actor-Critic (SAC) architecture, which serves as a common baseline (yu2020metaworld).
The multitask objective provides implicit regularization across multiple tasks trained together in a shared network (wuunderstanding), which can be formalized in a high-dimensional regression setting (yang2025precise). Training the model to learn representations useful across a diverse set of tasks helps prevent overfitting. A deeper analysis of the regularization effect behind multitask RL is an interesting question for future work.
@@
Another line of work involves data modeling, which aims to trace a model’s predictions back to its training data. One method is to retrain the model on a different subset of the dataset and estimate the Shapley value of each sample (ghorbani2019data). A commonly used technique is influence functions (koh2017understanding), which use Hessians to approximate the effect of removing a sample. Datamodels (ilyas2022datamodels) finds that a linear regression method can accurately approximate the outcome of deep nets trained with a subset of samples on computer vision tasks. li2023identification introduce a linear surrogate model for multitask learning and analyze the sample complexity of linear surrogate models. TRAK (park2023trak) further demonstrates that the approximate solution of linear regression computed with projected gradients delivers comparable results to the original linear model for ImageNet, CLIP, and BERT. Our work contributes to this literature by applying attribution methods to design RL algorithms.
Conclusion
@@ Acknowledgments
Clustering Algorithms
@@
Having introduced the notations, we start with the PAC-Bayes bound (mcallester2013pac), stated as follows.
@@
The proof will be deferred to Section Proof of Lemma 5. With these results ready, we will provide proof of the Hessian-based generalization bound.
@@
There exists an $\epsilon$-cover of $S$ with respect to the $\ell_{2}$-norm with at most $\max\Big(\big(\frac{3C}{\epsilon}\big)^{p},1\Big)$ elements; see, e.g., Example 5.8 (wainwright2019high). Let $T\subseteq S$ denote this cover. Recall that the Hessian $\nabla^{2}[\ell(f_{W}(x),y)]$ is $C_{1}$-Lipschitz for all $(W+U)\in S,W\in S$. Then we have
@@
By the matrix Bernstein inequality (vershynin2018high), we have
@@
The minimax optimal sample complexity to find an $\epsilon$-optimal $Q$-function in tabular $\gamma$-discounted MDPs is known to be $\tilde{O}\left({SA}{(1-\gamma)^{-3}\epsilon^{-2}}\right)$ (chi2025statistical). This rate is achieved by model-based algorithms. Standard synchronous $Q$-learning, however, is sub-optimal, requiring $\tilde{O}\left({SA}{(1-\gamma)^{-4}\epsilon^{-2}}\right)$ samples. This gap in horizon dependency was later closed by variance reduction techniques. For instance, a variance-reduced $Q$-learning algorithm (wainwright2019variance) achieves the minimax optimal rate, matching the lower bound up to logarithmic factors. See a recent monograph for extensive references (foster2023foundations). It is an interesting open question to study the sample complexity of $Q$-learning for multi-objective reinforcement learning. Another interesting question is to explore gradient estimation (li2025efficient; zhang2025linear) for designing ensemble methods in meta-RL. These are left for future work.
Omitted Experimentals
Environments. We summarize the RL environments: MT10, CartPole, Highway, LunarLander in Table 4. We adopt the Gymnasium library (towers2024gymnasium) for environment implementation.
Environment Goal State Action Reward

### v2

@@ Abstract
1 Introduction
Reinforcement learning (RL) is a technique for sequential decision-making in interactive environments, enabling agents to learn from feedback signals. A central challenge in RL is to develop agents that generalize across a wide range of tasks rather than solving a single task in isolation. For instance, a robot needs to master multiple related skills that share an underlying structure [yu2020metaworld]. Prior work has studied this problem for two closely related settings. The first, multitask reinforcement learning, seeks to find a single policy that performs well across a given set of objectives (or tasks) [sodhani2021multi, joshi2025benchmarking]. The second, meta-reinforcement learning, aims to learn a meta-initialization that can rapidly adapt to unseen tasks [finn2017model, rakelly2019efficient]. In both settings, however, the computational cost of evaluating and adapting a shared policy over many competing objectives grows quickly with $n$.
@@ 1 Introduction
This paper studies the problem of designing a (scalable) multi-objective optimizer for reinforcement learning, with extensions to policy meta-adaptation. Learning from a diverse set of tasks has been shown to improve generalization [thrun1998learning, cobbe2019quantifying, yin2019meta]. However, the main difficulty is modeling task relationships and how different tasks transfer information to one another when trained together in a complex network [wuunderstanding]. Existing approaches have considered measuring the similarity between gradient vectors across tasks to quantify task relationships [yu2020pcgrad]. Notably, this gradient-similarity-based measure corresponds to pairwise task-affinity scores, whereas in practice it is natural to model how multiple policies interact when trained together [li2023identification]. Exhaustively searching over all subsets of $\left\{1,2,\dots,n\right\}$ is computationally infeasible: evaluating each subset typically requires a full training procedure, and the number of subsets grows exponentially in $n$ [caruana1997multitask]. Greedy stepwise selection requires evaluation on $O(n^{2})$ subsets, which is again impractical for large $n$ [li2023boosting, li2024scalable].
@@
To overcome these challenges, we introduce a scalable algorithm for estimating policy performance on arbitrary task subsets without full training. Our algorithm starts by learning a single meta-policy across all tasks, then uses a first-order surrogate model—derived from a Taylor approximation around the meta-policy to efficiently estimate the outcome of fine-tuning on any subset. This approach is analogous to a random ensemble [zhang2025linear], which allows us to compute a pairwise task affinity matrix and partition the $n$ objectives into $k$ subgroups using a convex clustering procedure. The resulting groups can be trained separately using any multitask or meta-RL optimizers. This overall approach is illustrated in Figure 1.
@@
We validate our approach through extensive experiments on both Meta-World [yu2020metaworld] and robotic control benchmarks [towers2024gymnasium]. We find that the first-order approximation applied to the outputs of a policy network yields accurate estimates—typically within $2\%$ error margin—while being up to $26\times$ faster than full training. Our algorithm consistently outperforms multitask and meta-RL baselines [yang2020multi, sodhani2021multi, sun2022paco] by $16\%$ on average, and surpasses baseline grouping strategies (such as random grouping or gradient-similarity based grouping) by $23\%$ and $16\%$, respectively. These results highlight the effectiveness and scalability of our task-affinity estimation approach.
@@
Lastly, we conduct a theoretical analysis of the generalization errors of policy networks. The main technical approach is to compute (and measure) the trace of the Hessian of the loss surface [ju2022robust, ju2023generalization, zhangnoise]. See Theorem 1 for the precise statement. When evaluated on the above RL environments, we find that the Hessian-based generalization bounds match the scale of empirical generalization errors (see Figure 3).
@@ 2 Problem Setup
Pairwise task interactions can be characterized through this notion. For example, $R_{\{i,j\}}(\theta)$ captures the joint performance on tasks $T_{i}$ and $T_{j}$, providing a meaningful measure of their affinity. A key challenge in multitask learning is negative transfer, where conflicting gradients from two tasks degrade optimization efficiency and final performance [wuunderstanding]. This occurs when
@@
Our goal is to model the underlying relationship between the $n$ tasks, which can be used to group similar tasks or select a small representative subset for meta-learning. However, searching for the optimal partition requires evaluating $R_{S}(\theta)$ for multiple subsets $S\subseteq\mathcal{T}$, since the number of possible subsets is $2^{n}$. Greedy stepwise selection again requires evaluating $\mathcal{O}(n^{2})$ subsets, where each evaluation typically involves a full RL training or adaptation procedure [li2023boosting, li2024scalableft]. This is again intractable for large $n$.
@@ 3.1 Surrogate Modeling
We begin with the standard policy gradient objective used in algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal]:
@@
Random projection. Since the gradient vectors $g_{i,t}\in\mathbb{R}^{p}$ may be high-dimensional (with $p$ in the order of millions), we apply random projections to reduce the dimension down to a few hundred (in practice, $d=400$ suffices). This projection provably preserves the pairwise similarity between all pairs of gradient vectors via the Johnson–Lindenstrauss Lemma [johnson1984extensions]. We sample a random projection matrix $P\in\mathbb{R}^{p\times d}$, where $d\ll p$ and every entry of $P$ is independently drawn from a Gaussian $\mathcal{N}(0,d^{-1})$. We project the high-dimensional gradients to a $d$-dimensional space as
@@
3.2 Evaluation of First-Order Policy Approximation
@@
3.3 Task Affinity Grouping
@@
We then cluster tasks by solving a convex optimization problem that maximizes intra-cluster affinity, following a trace-regularized relaxation [awasthi2015relax]. This step is very fast since it runs on an $n$ by $n$ matrix with $n$ being at most several hundred, which can typically be solved in just a few seconds. Moreover, this procedure can be repeated with different values of $\lambda$ to determine the optimal number of clusters $k$ for downstream performance. For brevity, we defer a complete description of the convex relaxation program to the Appendix. The whole procedure is described in Algorithm 2.
@@
4 Experiments
@@
4.1 Experimental Setup
Environments. We evaluate our method on two types of benchmarks. First, we use MT10 from the Meta-World benchmark [yu2020metaworld], which consists of 10 diverse robotic manipulation tasks. While these tasks share common state and action spaces, their reward functions and dynamics vary (e.g., a goal position in one task may correspond to an object’s position in another). Second, we use three classic control environments from Gymnasium [towers2024gymnasium] and MO-Gymnasium [felten_toolkit_2023]: CartPole, Highway, and LunarLander. For each, we generate a distribution of 10 source tasks by altering key physical parameters (e.g., pole length for CartPole, traffic density for Highway, and gravity for LunarLander).
Baselines. We compare our method against two categories of baselines. (1) Multi-objective RL baselines: These represent standard and state-of-the-art multitask approaches for multitasks training: a single policy trained on all tasks; Soft Modularization [yang2020multi], which employs a routing-based selection mechanism; PaCo [sun2022paco], an ensemble method that composes policies from a shared parameter subspace; and CARE [sodhani2021multi], which performs inference using contextual information to adapt policies. (2) Grouping baselines: These methods use the same number of groups as our approach but employ different heuristic grouping procedures. This allows us to isolate the benefit of our affinity modeling. We include random grouping (where tasks are randomly assigned to $k$ groups) and gradient-similarity-based grouping (where tasks are clustered based on the cosine similarity of their respective policy gradients).
@@ 4.1 Experimental Setup
For the control environments, we form the meta-training set by randomly selecting one from each task group provided by PolicyGradEx. We then use MAML [finn2017model] to train a meta-policy on the assigned tasks.
@@
4.2 Results for Loss Function Estimation
@@
4.3 Results for Multi-Objective Optimization
Table 3: Comparison of our approach with several baseline methods. We report the average success rate on the Meta-World benchmark and the rewards (after adaptation) for three robotic control environments. As for the meta-RL setting, we several a subset of tasks for which we set as the source transfer tasks. We note that CARE does not apply to the control tasks in the meta-RL setting. We report the mean and standard deviation from five runs.
RL Environment Meta-World CartPole Highway LunarLander
Multi-task training [yu2020metaworld] $71.3_{\pm 1.2}\%$ $145.9_{\pm 9.0}$ $140.0_{\pm 4.6}$ $53.8_{\pm 14.6}$
Soft modularization [yang2020multi] $82.0_{\pm 1.1}\%$ $139.3_{\pm 9.5}$ $141.3_{\pm 3.5}$ $66.1_{\pm 13.0}$
PaCo [sun2022paco] $73.1_{\pm 1.1}\%$ $144.5_{\pm 5.2}$ $136.6_{\pm 6.4}$ $62.6_{\pm 11.4}$
CARE [sodhani2021multi] $84.0_{\pm 1.8}\%$ / / /
Gradient-similarity-based grouping [yu2020pcgrad] $69.6_{\pm 1.9}\%$ $142.0_{\pm 8.2}$ $135.6_{\pm 7.6}$ $80.8_{\pm 6.9}$
4.4 Generalization Error Measurements via Hessians
Next, we analyze the generalization behavior. by evaluating a sharpness measure based on the Hessian trace of the test loss for the policy network and the policy gradient loss. A smaller trace value suggests a flatter loss landscape. We implement Hutchinson’s estimator and leverage a faster version of this estimator [meyer2021hutch++] to estimate the Hessian trace of RL models. We compute the policy gradient loss from the target tasks and measure the difference between the training loss and test loss as the generalization error. Results for the largest eigenvalue of the Hessian are similar.
(a) Illustration of Hessian traces and generalization errors
@@ 4.5 Ablation Analysis
5 Generalization Error Analysis via Hessians
@@
The proof of Theorem 1 is based on a PAC-Bayes analysis [ju2022robust, zhangnoise], and the details are in the Appendix. In particular, a new aspect of this result is that we use an anisotropic perturbation in the prior and posterior distributions. We build on a linear PAC-Bayes bound (See Theorem 2 in the Appendix), and optimize the noise distribution to obtain the $\mathcal{H}$ measure.
@@
6 Related Work
Multi-objective optimization for reinforcement learning. Early works on multitask learning use hard parameter sharing, where a single neural network backbone is shared among all tasks, with only the final layers or heads being task-specific [yu2020metaworld]. A canonical example is the multi-head Soft Actor-Critic (SAC) architecture, which serves as a common baseline [yu2020metaworld].
The multitask objective provides implicit regularization across multiple tasks trained together in a shared network [wuunderstanding], which can be formalized in a high-dimensional regression setting [yang2025precise]. Training the model to learn representations useful across a diverse set of tasks helps prevent overfitting. A deeper analysis of the regularization effect behind multitask RL is an interesting question for future work.
@@ 6 Related Work
Another line of work involves data modeling, which aims to trace a model’s predictions back to its training data. One method is to retrain the model on a different subset of the dataset and estimate the Shapley value of each sample [ghorbani2019data]. A commonly used technique is influence functions [koh2017understanding], which use Hessians to approximate the effect of removing a sample. Datamodels [ilyas2022datamodels] finds that a linear regression method can accurately approximate the outcome of deep nets trained with a subset of samples on computer vision tasks. [li2023identification] introduce a linear surrogate model for multitask learning and analyze the sample complexity of linear surrogate models. TRAK [park2023trak] further demonstrates that the approximate solution of linear regression computed with projected gradients delivers comparable results to the original linear model for ImageNet, CLIP, and BERT. Our work contributes to this literature by applying attribution methods to design RL algorithms.
7 Conclusion
@@ Acknowledgments
{refcontext}
[sorting=nyt]
Appendix A Clustering Algorithms
@@ Appendix B Proof of Theorem 1
Having introduced the notations, we start with the PAC-Bayes bound [mcallester2013pac], stated as follows.
@@
The proof will be deferred to Section B.0.1. With these results ready, we will provide proof of the Hessian-based generalization bound.
@@ B.0.1 Proof of Lemma 6
There exists an $\epsilon$-cover of $S$ with respect to the $\ell_{2}$-norm with at most $\max\Big(\big(\frac{3C}{\epsilon}\big)^{p},1\Big)$ elements; see, e.g., Example 5.8 [wainwright2019high]. Let $T\subseteq S$ denote this cover. Recall that the Hessian $\nabla^{2}[\ell(f_{W}(x),y)]$ is $C_{1}$-Lipschitz for all $(W+U)\in S,W\in S$. Then we have
@@
By the matrix Bernstein inequality [vershynin2018high], we have
@@
The minimax optimal sample complexity to find an $\epsilon$-optimal $Q$-function in tabular $\gamma$-discounted MDPs is known to be $\tilde{O}\left({SA}{(1-\gamma)^{-3}\epsilon^{-2}}\right)$ [chi2025statistical]. This rate is achieved by model-based algorithms. Standard synchronous $Q$-learning, however, is sub-optimal, requiring $\tilde{O}\left({SA}{(1-\gamma)^{-4}\epsilon^{-2}}\right)$ samples. This gap in horizon dependency was later closed by variance reduction techniques. For instance, a variance-reduced $Q$-learning algorithm [wainwright2019variance] achieves the minimax optimal rate, matching the lower bound up to logarithmic factors. See a recent monograph for extensive references [foster2023foundations]. It is an interesting open question to study the sample complexity of $Q$-learning for multi-objective reinforcement learning. Another interesting question is to explore gradient estimation [li2025efficient, zhang2025linear] for designing ensemble methods in meta-RL. These are left for future work.
Appendix C Omitted Experimentals
Environments. We summarize the RL environments: MT10, CartPole, Highway, LunarLander in Table 4. We adopt the Gymnasium library [towers2024gymnasium] for environment implementation.
Table 4: A summary of RL environments tested in our experiments.
Env. Goal State Action Reward